OpenStudy (anonymous):

What is the multiplicative rate of change for the exponential function f(x) = 2(5/2)–x(to the negative X)? A.0.4 B.0.6 C.1.5 D.2.5

OpenStudy (perl):

Is this the function? $$\Large f(x) = 2\left(\frac52 \right) ^{-x} $$

OpenStudy (anonymous):

yeah

OpenStudy (perl):

first lets make the exponent positive

OpenStudy (perl):

$$\Large f(x) = 2\left(\frac52 \right) ^{-x} = 2 \left(\frac25 \right) ^{x} $$ 2/5 is the multiplicative rate of change, I think.

OpenStudy (perl):

since 2 is being multiplied by 2/5, then (2/5)^2 , etc

OpenStudy (perl):

what does 2/5 equal to in decimal?

OpenStudy (anonymous):

0.4